# NTPP: Generative Speech Language Modeling for Dual-Channel Spoken Dialogue via Next-Token-Pair Prediction

### v2

@@ 5.1 Dataset and Baselines
Baselines We evaluate NTPP’s performance by comparing it with established generative speech systems. In terms of turn-taking statistics within generated dialogues, we compared NTPP specifically with dGSLM (Nguyen et al., 2023), as it represents a comparable full-duplex generative model trained on the Fisher dataset. The evaluation in (Nguyen et al., 2023) involved 50 prompts, each generating 50 continuations, with reported results for both their dialogue model and a cascaded topline model, which consists of an Automatic Speech Recognition(ASR) model, followed by a text-based language model and a Text-To-Speech (TTS) module. Following the settings of (Nguyen et al., 2023), we select wav2vec2-large(Baevski et al., 2020), KenLM(Heafield, 2011), and Google TTS API as the modules respectively. For assessing the meaningfulness and naturalness of the interaction, we extend our comparison to include Moshi (Défossez et al., 2024) and SyncLLM (Pope et al., 2023a). Notably, SyncLLM is a full-duplex dialogue agent designed for real-time, overlapping speech interactions through the joint, streaming processing of speech input and output.
@@
Table 2: Linguistic quality and turn-taking statistics of generated dialogues, including the number of turn-taking events and cumulative durations per minute, compared to the ground truth. For each prompt of NTPP, we generated 32 continuations across three different temperature settings [0.1, 0.5, 0.9], as temperature significantly impacts the results.
@@ 5.2 SLM Training &\&& Implementation Details
LLM Foundation Models.
To obtain SLMs, we begin with three popular LLM backbones: LLaMA 3.1–8B (Dubey et al., 2024), Mistral-7B-v0.1 (Jiang et al., 2023a), and Gemma-2-9B (Team et al., 2024). We evaluate the performance of SLMs obtained by training these LLMs with the NTP objective on a large-scale single-channel audio dataset. Perplexity on the test set is used as the evaluation metric. Figure 5 shows the training loss curves over time for all three models. Each model demonstrates a consistent downward trend, indicating effective learning. Mistral-7B and Gemma show comparable learning dynamics, while LLaMA 3.1—known for its strong reasoning capabilities in text—achieves lower training loss more quickly. This finding supports our hypothesis that stronger text-based models serve as more effective initializations for continual speech learning, consistent with the perspective of treating audio as a new language.
Perplexity curves reveal that NTPP, when trained solely on audio data (“w/o Text”), exhibits significantly faster convergence and consistently achieves lower perplexity throughout the training process compared to the model trained with both audio and ASR text data (“w Text”). The “w/o Text” approach signifies training exclusively with audio. In contrast, the “w Text” approach integrates an ASR task, using both audio and its corresponding text data during training. The results show that the audio-only setting leads to significantly faster convergence and consistently lower perplexity throughout training. This indicates that eliminating the ASR transcript—which may introduce recognition errors or redundancy—allows the model to focus on more informative acoustic cues, thereby facilitating more stable and efficient learning. In contrast, the inclusion of textual supervision appears to slow optimization and result in higher perplexity, suggesting potential modality interference.
@@ 5.3 Turn-taking Event Distribution
Following the experimental setting in dGSLM (Nguyen et al., 2023), we evaluate the dialogue systems with turn-taking capabilities using corpus-level statistics (Ardila et al., 2019) and testing on Fisher dataset (Cieri et al., 2004). We evaluate the linguistic quality and turn-taking dynamics of generated dialogues using various models, as shown in Table 2. The detailed evaluation settings are in the Appendix 2. LSLM (Ma et al., 2024) integrates speaker channels at the embedding layer and separates them in the final layer, demonstrates a notable reduction in the number of Inter-Pausal Units (IPUs) and gaps, indicating smoother transitions between speakers. The dGSLM (Nguyen et al., 2023), particularly with the cross-attention module, shows a significant decrease in the cumulative duration of pauses and gaps, suggesting more fluid and continuous dialogue. Comparatively, NTPP exhibit balanced performance with moderate reductions in both the number and duration of turn-taking events, highlighting their potential for generating natural and coherent dialogues. These findings underscore the importance of model architecture in optimizing dialogue flow and linguistic quality.
@@ 5.4 Interruptions & Reflective Pause
We further develop a comprehensive evaluation framework comprising 400 diverse conversational scenarios specifically designed to systematically capture natural dialogue dynamics, with an emphasis on pauses and interruptions. These scenarios are carefully crafted using GPT-4 to reflect the complexity and nuance of human interactions. We then use ChatTTS(Fathullah et al., 2024) to synthesize high-quality speech from the generated text, effectively mimicking the acoustic characteristics of real-world conversations. As shown in Figure 6, NTPP demonstrates closer alignment with human reference judgments in both speaking-up and interruption scenarios when acting as a listener. This suggests that NTPP more effectively captures the subtleties of human conversational behavior compared to other models.
@@
Table 3: Meaningfulness (Meaning.) and Naturalness (Nat.) (scores 1-5) mean estimates and standard errors (in parentheses), aggregated overall and for Fisher and CANDOR subsets.
@@ 5.5 Human Evaluation
We follow the evaluation protocol from (Veluri et al., 2024) and conduct a human study involving 25 annotators with native-level English proficiency. Adopting the Mean Opinion Score (MOS) framework, we use a 5-point Likert scale to evaluate the Naturalness (N-MOS) of turn-taking and the Meaningfulness (M-MOS) of the generated dialogue content. Table 3 presents a comparison of NTPP with various baselines in terms of both naturalness and meaningfulness. Notably, dGSLM and SyncLLM are trained solely on the Fisher dataset—the only real-world spoken dialogue dataset used in their training. We also include performance comparisons on the out-of-distribution Candor test set (Reece et al., 2023) to assess generalization.
@@ 5.6 Inference Latency
In multi-turn interactions, latency is a crucial metric for assessing the performance of speech interaction models, as it captures the time between receiving a speech input and initiating a response—directly impacting user experience. As shown in Figure 7, our NTPP consistently achieves lower inference latency than Moshi, especially as the number of turn-taking rounds increases. We attribute this advantage to NTPP’s efficient memory usage: it maintains a single KV Cache, while Moshi requires two separate KV Caches for the two speaker channels. This difference becomes increasingly significant as conversational context grows longer.
@@ 5.7 Speaker Independence
To assess the speaker-independence of various models, we conduct a speaker-swapped evaluation. All models are first trained on the dual-channel Fisher dataset (Cieri et al., 2004) using the canonical speaker order. During inference, we reverse the input speaker sequence without any additional fine-tuning or adaptation. This setup allows us to test whether a model’s performance remains stable when the roles of the two speakers are exchanged—an essential indicator of robustness and generalization in dialogue turn-taking modeling. We evaluate this on both the training and test sets. As shown in Table 4, both dGSLM and NTPP exhibit minimal variation in key turn-taking metrics under speaker-swapped conditions—nearly zero variation on the training set and consistently low variation on the test set—demonstrating strong speaker-independent behavior. In contrast, Moshi shows substantial deviations in metrics such as the number and duration of IPUs, pauses, gaps, and overlaps. These findings suggest that Moshi relies on speaker-conditioned generation (due to its modeling of conditional distributions), resulting in degraded performance when the input speaker order is reversed.
Figure 8: Comparative performance analysis. (a) Ablation study of NTPP model stages, illustrating perplexity versus training steps for the full two-stage model and variants lacking either stage one or stage two. (b) Comparison of VQ and RVQ model training loss as a function of training steps.
@@ 5.8 Ablation Studies
We further investigate the importance of two-stage training and the use of an RVQ tokenizer in this section. As shown in Figure 8, our two-stage training strategy consistently achieves lower perplexity compared to the one-stage approach, which omits pretraining on single-channel audio. This suggests that pretraining on single-channel audio provides a strong foundation, significantly improving performance on subsequent dual-channel speech learning. As expected, omitting the second-stage NTPP training on dual-channel speech also leads to performance degradation. Figure 8 compares training loss across different tokenizers, with RVQ yielding consistently lower loss, highlighting the importance of developing extended solutions tailored to this tokenizer.
@@ C.1 Hyper-parameter Settings
C.2 Dialogue linguistic quality

### v4

@@ 5.1 Dataset and Baselines
Baselines We evaluate NTPP’s performance by comparing it with established generative speech systems. In terms of turn-taking statistics within generated dialogues, we compared NTPP specifically with dGSLM (Nguyen et al., 2023), as it represents a comparable full-duplex generative model trained on the Fisher dataset. Following dGSLM’s setting (Nguyen et al., 2023), we include a cascaded baseline which consists of an Automatic Speech Recognition(ASR) model, followed by a text-based language model and a Text-To-Speech (TTS) module. Following the settings of (Nguyen et al., 2023), we select wav2vec2-large(Baevski et al., 2020), KenLM(Heafield, 2011), and Google TTS API as the modules respectively. For assessing the meaningfulness and naturalness of the interaction, we extend our comparison to include Moshi (Défossez et al., 2024) and SyncLLM (Pope et al., 2023a). Notably, SyncLLM is a full-duplex dialogue agent designed for real-time, overlapping speech interactions through the joint, streaming processing of speech input and output.
@@
Table 2: Turn-taking statistics (event occurrences and durations/min) for generated dialogues, compared to ground truth using mean absolute differences ($|\Delta|$) between predicted and ground truth values. NTPP model results cover different temperature settings [0.1, 0.5, 0.9]. Lower $|\Delta|$ values indicate better performance, with the best results highlighted in bold.
@@ 5.2 SLM Training &\&& Implementation Details
LLM Backbones.
We leverage three well-known LLMs as the backbone of our SLM: LLaMA 3.1–8B (Dubey et al., 2024), Mistral-7B-v0.1 (Jiang et al., 2023a), and Gemma-2-9B (Team et al., 2024). We evaluate the performance of SLMs obtained by training these LLMs with the NTP objective on a large-scale single-channel audio dataset. Perplexity on the test set is used as the evaluation metric. Figure 5 shows the training loss curves over time for all three models. Each model demonstrates a consistent downward trend, indicating effective learning. Mistral-7B and Gemma show comparable learning dynamics, while LLaMA 3.1—known for its strong reasoning capabilities in text—achieves lower training loss more quickly. This finding supports our hypothesis that stronger text-based models serve as more effective initializations for continual speech learning, consistent with the perspective of treating audio as a new language.
Perplexity curves highlights that during the pre-training phase, the NTPP model demonstrates superior performance when trained exclusively on audio data (“w/o Text”) compared to when trained with both audio and its ASR text transcriptions (“w Text”). Specifically, the audio-only approach leads to significantly faster convergence and consistently lower perplexity—a measure indicating better predictive capability of the model. The results show that the audio-only setting leads to significantly faster convergence and consistently lower perplexity throughout training. This indicates that eliminating the ASR transcript—which may introduce recognition errors or redundancy—allows the model to focus on more informative acoustic cues, thereby facilitating more stable and efficient learning. In contrast, the inclusion of textual supervision appears to slow optimization and result in higher perplexity, suggesting potential modality interference.
@@ 5.3 Turn-taking Event Distribution
Following the experimental setting in dGSLM (Nguyen et al., 2023), we evaluate the dialogue systems with turn-taking capabilities using corpus-level statistics (Ardila et al., 2019) and testing on the Fisher dataset (Cieri et al., 2004). We evaluate the linguistic quality and turn-taking dynamics of generated dialogues using various models, as shown in Table 2. The detailed evaluation settings are in the Appendix C.2. LSLM (Ma et al., 2024) integrates speaker channels at the embedding layer and only predicts the speech tokens from the assistant channel, demonstrating a notable reduction in the number of Inter-Pausal Units (IPUs) and gaps, indicating smoother transitions between speakers. The dGSLM (Nguyen et al., 2023), particularly with the cross-attention module, shows a significant decrease in the cumulative duration of pauses and gaps, suggesting more fluid and continuous dialogue. Comparatively, NTPP exhibits balanced performance with moderate reductions in both the number and duration of turn-taking events, highlighting their potential for generating natural and coherent dialogues. These findings underscore the importance of model architecture in optimizing dialogue flow.
@@ 5.4 Interruptions & Reflective Pause
We further develop a comprehensive evaluation framework comprising 400 diverse conversational scenarios specifically designed to systematically capture natural dialogue dynamics, with an emphasis on pauses and interruptions. These scenarios are carefully crafted using GPT-4 to reflect the complexity and nuance of human interactions. We then use ChatTTS(cha, 2024) to synthesize high-quality speech from the generated text, effectively mimicking the acoustic characteristics of real-world conversations. As shown in Figure 6, NTPP demonstrates closer alignment with human reference judgments, which are meticulously established through manual annotation by human evaluators, in both speaking-up and interruption scenarios when acting as a listener. For instance, informing its response strategy, NTPP employs Voice Activity Detection (VAD) and considers a silence state to have been reached after 200ms of continuous non-speech is detected, facilitating its decisions on when to speak or acknowledge a pause. This suggests that NTPP more effectively captures the subtleties of human conversational behavior compared to other models.
@@
Table 3: Human evaluation results across different SLMs. Meaningfulness (Meaning.) and Naturalness (Nat.) scores (ranging from 1 to 5) represent mean estimates and their standard errors (shown in parentheses), reported both overall and separately for the Fisher and CANDOR datasets.
@@ 5.5 Human Evaluation
We follow the evaluation protocol from (Veluri et al., 2024) and conduct a human study involving 25 annotators with native-level English proficiency. Adopting the Mean Opinion Score (MOS) framework, we use a 5-point Likert scale to evaluate the Naturalness (N-MOS) of turn-taking and the Meaningfulness (M-MOS) of the generated dialogue content. Table 3 presents a comparison of NTPP with various baselines in terms of both naturalness and meaningfulness. We also include performance comparisons on the out-of-distribution Candor test set (Reece et al., 2023) to assess generalization.
@@ 5.6 Inference Latency
In multi-turn interactions, latency is a crucial metric for assessing the performance of speech interaction models, as it captures the time between receiving the end of input speech and the start of output speech, directly impacting user experience. As shown in Figure 7, our NTPP consistently achieves lower inference latency than Moshi, especially as the number of turn-taking rounds increases. We attribute this advantage to NTPP’s efficient memory usage: it maintains a single KV Cache, while Moshi requires two separate KV Caches for the two separate transformers. This difference becomes increasingly significant as conversational context grows longer.
@@ 5.7 Speaker Independence
To assess the speaker-independence of various models, we conduct a speaker-swapped evaluation. All models are first trained on the dual-channel Fisher dataset (Cieri et al., 2004) using the canonical speaker order. We reverse the input speaker sequence without any additional fine-tuning or adaptation. This setup allows us to test whether a model’s performance remains stable when the roles of the two speakers are exchanged—an essential indicator of robustness and generalization in dialogue turn-taking modeling. We evaluate this on both the training and test sets. As shown in Table 4, both dGSLM and NTPP exhibit minimal variation in key turn-taking metrics under speaker-swapped conditions—nearly zero variation on the training set and consistently low variation on the test set—demonstrating strong speaker-independent behavior. Moshi shows substantial deviations in metrics such as the number and duration of IPUs, pauses, gaps, and overlaps. These findings suggest that Moshi relies on speaker-conditioned generation (due to its modeling of conditional distributions), resulting in degraded performance when the input speaker order is reversed.
Figure 8: Comparative performance analysis. (a) Ablation study of NTPP model training stages, illustrating perplexity versus training steps for the full two-stage model and variants lacking either stage one or stage two. For better visualization, we trim the stage one training steps to be the same as stage two. (b) Comparison of VQ and RVQ model training loss as a function of training steps.
@@ 5.8 Ablation Studies
We further investigate the importance of two-stage training and the use of an RVQ tokenizer in this section. As shown in Figure 8 (a), our two-stage training strategy consistently achieves lower perplexity compared to the one-stage approach, which omits pretraining on single-channel audio. This suggests that pretraining on single-channel audio provides a strong foundation, significantly improving performance on subsequent dual-channel speech learning. As expected, omitting the second-stage NTPP training on dual-channel speech also leads to performance degradation. Figure 8 (b) compares training loss between VQ and RVQ tokenizers, with RVQ yielding consistently lower loss, highlighting the importance of developing extended solutions tailored to this tokenizer.
@@ C.1 Hyper-parameter Settings
C.2 Dialogue turn-taking statistics evaluation